OpenStudy (dangerousjesse):

I just bought a 24 pack of cinnamon flavored Trident gum for $12.67, and each packet of gum contains 18 pieces. A few weeks ago, I bought a small pack of cinnamon Tooty Fruity gum for $0.25 and it contained 9 pieces of gum. Which one is the better buy? (Real-life situation, I found the answer earlier. Just want to see how long it takes.

OpenStudy (yanasidlinskiy):

Tooty Fruity? Hahha!!! I get the Minty Kind!!xD!! No offense!! Though..xD!! Give me some, btw!!

OpenStudy (x3_drummerchick):

wouldnt it be 12.67 / .25 which equals roughly 50 pieces of gum for a quarter a pack. 18 * 24 = 432 pieces

OpenStudy (x3_drummerchick):

so the 24 pack is better

OpenStudy (aaronq):

the price in bulk is \(\dfrac{$12.67}{24*18~pieces}=$0.029328/piece\approx $0.0294/piece\) the price in smaller amounts is \(\dfrac{$0.25}{8~pieces}=$0.03125/piece \approx $0.0313 /piece\) you're basically saving 2 cents per piece of gum.

OpenStudy (aaronq):

ah crap it was \(\dfrac{$0.25}{9~pieces}=$0.027777 /piece \approx $0.0278/piece\) so the smaller pack of 9 pieces is cheaper by about 2 cents

OpenStudy (dangerousjesse):

The easier way to do it @aaronq .25\(\times\)2=.50 .50\(\times\)24=12 12 dollars for a 24 pack of 18-piece-per-packet tooty fruity 12.67 dollars for a 24 pack of 18-piece-per-packet trident

OpenStudy (aaronq):

true true, it's pretty easy either way